OpenStudy (anonymous):

8 + (-13/5) + (4/3)=

OpenStudy (anonymous):

\[8 + (-\frac{ 13 }{ 5 }) + (\frac{ 4 }{ 3 })=\]

OpenStudy (anonymous):

get the LCD.. :)

OpenStudy (anonymous):

How do you do that with a problem like this?

OpenStudy (anonymous):

haha.. i don't know .. but I know how to solve. LOL

OpenStudy (anonymous):

Can you show me how to solve it then, please?

OpenStudy (anonymous):

do you know what is LCD?

OpenStudy (anonymous):

Least common difference?

OpenStudy (anonymous):

no.. least common denominator..

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

so, do you know that??

OpenStudy (anonymous):

I think I do. It has been awhile since I've done it though.

OpenStudy (anonymous):

okay.. so get the LCD. try it first on your own. :)

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

yes.. :) then divide and multiply it to the numbers..

OpenStudy (anonymous):

8+(-39/15) + (20/15)

OpenStudy (anonymous):

no.. wrong. :) include 8 in dividing and multiplying.. :)

OpenStudy (anonymous):

101/15

OpenStudy (anonymous):

yes.. :) you're right.. :) you can translate that in mixed fractions or in decimal form, if you don't want the improper fraction to be your final answer,.. :)

OpenStudy (anonymous):

Would I divide 15 into 101 to make it a mixed fraction?

OpenStudy (anonymous):

\[8 + \frac{(-13)}{5} +\frac{ 4}{3}=8 - \frac{13}{5} +\frac{ 4}{3}=8 +\frac{ 4}{3}- \frac{13}{5} \] \[=8 +\frac{ 4 \times 5 - 13 \times 3}{3 \times 5} =8 +\frac{ 20 - 39}{15}=8 +\frac{ - 19}{15}\] \[=\frac{ 8 \times 15- 19}{15}=\frac{ 120- 19}{15}=\frac{ 101}{15}\]

OpenStudy (anonymous):

6 11/15

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yes.. :) am, welcome. :)